9. Calculate the pressure exerted by 5.00 mol of Co2 in a 1.00 L vessel at 300 K (a) assuming the gas behaves ideally and (b) using the van der Waals equation. (a= 3.610 Latm/mol, and b = 0.0429 L/mol).
Answer #1

9)
a)
P*V = n*R*T
P*1 = 5.00*0.082*300
P = 123 atm

b)
(P+an^2/V^2)*(V-nb) = n*R*T
(P+3.610*25)*(1-5*0.0429) = 5.00*8.314*300
(P+90.25)*0.786 = 123
(P+90.25) = 156.4
P = 66.15 atm
= 66.2 atm